the midpoint of the line segment from P1 to P2 is(-1,7). If P1=(-7,8) what is P2
Answer by lwsshak3(11628)

the midpoint of the line segment from P1 to P2 is(-1,7). If P1=(-7,8) what is P2
**
midpoint formula: (x1+x2)/2,(y1+y2)/2
For given problem:
P1=(-7,8)
P2=(x,y)
Midpoint=(-1,7)
..
(x-7)/2, (y+8)/2=(-1,7)
(x-7)/2=-1
x-7=-2
x=5
..
(y+8)/2=7
y+8=14
y=6
..
P2=(5,6)
check: (5-7)/2,(6+8)/2=(-1,7)